The Jolly Company sells candy and issued an RFQ to purchase a line of chocolates that are made do their specifications. After inspecting the samples for quality, the following results were determined for the three suppliers (Supplier X, Supplier Y, and Supplier Z), as presented in Table 1 (below).

Quality rating were determined by the look and taste of the samples. The quality of the samples from Supplier X, the low-cost supplier, did not taste and look nearly as good as those from Supplier Y and Z. The prices quoted per pound were as follows: Supplier X is $8.40, Supplier Y is $9.00, and Supplier Z is $9.40 per pound.

Table 1: Results for Three Suppliers (X, Y, and Z)

Supplier X

Supplier Y

Supplier Z

Criteria

Weight

Rating

Score

Rating

Score

Rating

Score

Price

Quality

Delivery Reliability

Customer Service

60%

20

10

10

5

2

3

2

4

5

4

5

3

5

3

4

Total Score

100%

Inquiries

  1. Showing all work, calculate the Total Score for each of the three (3) suppliers: Supplier X, Supplier Y, and Supplier Z.
  2. Which supplier should be selected based on the scores?
  3. Explain the underlying reason the lowest price supplier was not chosen.
Answer #1

1)

Total weighted score for supplier X = .6*5+.2*2+.1*3+.1*2 = 3.9

Total weighted score for supplier Y = .6*4+.2*5+.1*4+.1*5 = 4.3

Total weighted score for supplier Z = .6*3+.2*5+.1*3+.1*4 = 3.5

2)

Total weighted score of supplier Y is the highest.

Therefore, based on the scores, supplier Y should be selected.

3)

Supplier X has quoted the lowest price. However, it is still not a recommended choice, because its rating in terms of quality, customer service and delivery reliability are low compared to that of supplier Y. Therefore, considering the overall impact of all the four factors according to their respective weightages, supplier Y should be selected instead of X